OpenStudy (anonymous):

(6(x+h)^2-2(x+h)+7-(6x^2-2x+7))/h Evaluate Lim h -->0

OpenStudy (mimi_x3):

6(x+h)^2 - 2(x+h) + 7 - (6x^2-2x+7) 6(h^2+2hx+x^2)-2x-2h+7-6x^2+2x-7 / h 6h^2 + 12hx + 6x^2 - 2x - 2h + 7 - 6x^2 + 2x - 7 / h 12hx -2h / h h(12x-2) / h = 12x-2 Like that ?

OpenStudy (anonymous):

From step 3 to 4, I ended up with 6h+12x-2.Could you please explain?

OpenStudy (mimi_x3):

Woops, I made a stupid mistake sorry. You're right it's 6h+12x-2 Want me to do it again ?

OpenStudy (anonymous):

No no, its fine. Could you please continue from this step though? Thanks for your help!

OpenStudy (mimi_x3):

h-> 0 , 6h + 12x - 2 = 6(0) + 12x -2 = 12x-2 It's right now dw http://www.wolframalpha.com/input/?i=limit+h+tend+to+0%286%28x%2Bh%29%5E2-2%28x%2Bh%29%2B7-%286x%5E2-2x%2B7%29%29%2Fh

OpenStudy (anonymous):

Thanks!

OpenStudy (mimi_x3):

yw (: